# Analytical Insights into Constant-Roll Condition: Extending the Paradigm to Non-Canonical Models

### v1

@@ 5.3 PBH Formation
$z^{\prime\prime}/(a^{2}H^{2}z)$
$N$$N$$N$$N$$N$$Log(k)$CanonicFCRHCRPCR(a)(b)(c)(d)(e)(f)
Figure 10: (a) constant-roll parameter $\beta$, (b) power spectrum $\mathcal{P}_{s}$, (c) first slow-roll parameter $\epsilon_{1}$, (d) second slow-roll parameter $\epsilon_{2}$, (e) parameter $z^{\prime}/(aHz)$, and (f) effective potential $z^{\prime\prime}/(a^{2}H^{2}z)$ for the constant $k$ inflation. In all panels except panel (a), green, blue, and yellow lines represent FCR, HCR, and PCR conditions, respectively. Dashed gray lines correspond to the canonical case. There are four constant-roll stages with $\beta(N)\approx-0.017$, $-1.45$, $0.018$, and $5$ separated by three vertical grid lines at $N=10$, $20.5$, and $60$ e-folds. The initial stage of the scenario is a slow-roll case that meets the CMB constraints at the pivot point $k_{\text{cmb}}=0.05\text{Mpc}^{-1}$. In the second stage the constant-roll parameter is chosen smaller than $-1$ in order to induce the formation of PBHs, but larger than $-3/2$ to preserve the attractor behavior. Upon reaching the desired amplitude of power spectrum, we enter another slow-roll stage to prevent overproduction of PBHs. During the final stage of the scenario, the constant-roll parameter is large and positive, resulting in a quick exit from inflation.
We start by constructing a multiple-stage constant-roll parameter as follows:{sequation}[] β(N) =-49310 (20001953+tanh(3 (N-15))) (3940-tanh(3 (N-20.5))) (137136 + tanh(3 (N-60))). This parameter (plotted in panel (a) of Figure 10) describes four stages of constant-roll inflation (with $\beta=-0.017$, $-1.45$, $0.018$, and $5$, respectively). The stages are connected using proper smooth functions. The constant-roll parameter is set to $-0.017$ in the first stage in order to comply with Planck data (see discussion after equation (2.36)). At this stage we normalized the power spectrum to meet the CMB constraints [74], $\mathcal{P}_{s}=2.1\times 10^{-9}$ at $k_{\text{cmb}}=0.05\text{Mpc}^{-1}$. After 15 e-folds we switch to a large $\beta$ stage responsible for enhancing the power spectrum. We take the constant-roll parameter smaller than $-1$ to have negative $z^{\prime}/z$, but larger than $-3/2$ in order to preserve the attractor behavior of the model. At $21.5$ e-folds, when the power has increased around $10^{7}$ times, we end this stage. The third stage is another slow-roll stage with $\beta=0.018$ that is compatible with Planck data (see discussion after equation (2.35)). At the end, we enter a short stage with large positive $\beta$ at 60 e-folds, that leads to a sharp increase in $\dot{\phi}$ and accordingly $\epsilon_{1}$, allowing us to reach $\epsilon_{1}=1$ and exit inflation gracefully.
In the canonical case, we use parameter (5.3) in equation (2.6) and solve the equations of motion and Mukhanov-Sasaki equation (A.2) numerically to derive power spectrum and dynamical parameters. The results are illustrated in Figure 10 (dashed gray lines). As can be seen, the power has increased by about $10^{7}$ times, which is appropriate for PBH formation as suggested in references [57, 58]. In this scenario, when $\beta$ changes, we might automatically switch to another constant-roll solution. Solutions (2.15) and (2.13) can be used to describe the first and second stages respectively, and solution (2.14) can be used to describe the last two stages.
For $k$ inflation the situation is more complicated. In this model we start inflation in $\gamma\dot{\phi}^{2}\gg 1$ phase when (5.9) is valid. In the first stage, since $\beta$ is small and $\dot{\phi}$ decays slowly, there is not enough time to transit to the canonical case. During the second stage, $|\beta|$ increases significantly leading to a quick decay in $\dot{\phi}$ and the transition to the canonical case. During the third stage $\beta$ becomes positive and $\dot{\phi}$ begins to grow. However, since $\beta$ is small we remain in the canonical case throughout. In the last stage, as $\beta$ is large and positive, $\dot{\phi}$ grows dramatically fast and we transit back to $\gamma\dot{\phi}^{2}\gg 1$ limit before exiting inflation.
After using (5.3) in three constant-roll conditions (5.6), (5.7), and (5.8), we numerically solved dynamical equations to derive the power spectrum and dynamical parameters. Figure 10 shows the results (the green, blue, and yellow lines represent FCR, HCR, and PCR conditions, respectively). It should be noted that here we want to compare three constant-roll conditions within a single scenario. We are concerned about the effect of changing the constant-roll condition on the model, not the compatibility of each power spectrum with the data.
Looking at panels (c) and (d) of Figure 10, it is evident that the slow-roll parameters for the HCR condition are identical to those for the canonical case. It is due to the fact that for the HCR condition slow-roll parameters can be derived irrespective of action. However, since the sound speed is not constant, the power spectrum will be different. As can be seen in panels (e) and (f) of Figure 10, parameters $z^{\prime}/(aHz)$ and $z^{\prime\prime}/(a^{2}H^{2}z)$ for the HCR condition differ slightly from the canonical case, which is the direct result of varying sound speed (also see equations (5.27) and (A.3)).
Panel (c) of Figure 10 also shows that in all stages of the scenario except the third stage, the slope of $Log(\epsilon_{1})$ differs for three constant-roll conditions. This is because at third stage we reach the canonical case regardless of our choice of constant-roll condition. This point becomes more clear when we look at the second slow-roll parameter in panel (d) of Figure 10. Despite reaching a single path at about 20 e-folds, $\epsilon_{2}$ functions are different in earlier times when we have not reached the canonical limit yet. In early times of inflation when approximation (5.9) is valid, in order to get similar results from three constant-roll conditions we must choose $6\beta_{\text{\tiny{F}}}=3\beta_{\text{\tiny{H}}}=2\beta_{\text{\tiny{V}}}$ (see equation (5.10)), but our choice is $\beta_{\text{\tiny{F}}}=\beta_{\text{\tiny{H}}}=\beta_{\text{\tiny{V}}}=\beta(N)$, where $\beta(N)$ is defined in (5.3). However, when we enter the canonical phase, our choice $\beta_{\text{\tiny{F}}}=\beta_{\text{\tiny{H}}}=\beta_{\text{\tiny{V}}}=\beta(N)$ leads to similar results for three constant-roll conditions (see discussion after equation (5.9)). Therefore, we can expect a similar dynamic for three constant-roll conditions at third stage since we are in the canonical limit.
@@
Looking at panel (e) of Figure 10, we see that in all cases $z^{\prime}/(aHz)$ transiently changes sign, which is necessary to enhance the power spectrum. The more it goes toward negative signs the stronger will be the driving force that enhances the power. Therefore, by looking at $z^{\prime}/(aHz)$ we expect that the FCR condition, followed by the HCR and PCR conditions, will result in greater power spectrum enhancement. Which is the case as can be seen in panel (b) of Figure 10.
@@ 6 Summary and Conclusion
This study explored three methods for imposing the constant-roll condition on non-canonical models. We used a general action as a toy model to introduce three distinct constant-roll conditions, which are all equivalent in the canonical limit but may lead to different results in non-canonical cases. In the first approach, in which a more conventional form of the constant-roll condition in the literature is used, the condition was defined to ensure that the inflaton rolling rate remains constant throughout the inflation. Imposing another type of the constant-roll condition ensures a constant $\epsilon_{2}$ in the model. This second condition enabled us to derive the time profile of the Hubble and slow-roll parameters analytically, even without knowing the exact form of the action. For models with constant sound speed the power spectrum and spectral index of the model were also derived. In these two approaches, we may not obtain a flat USR potential for any value of the constant-roll parameter. Therefore, in our third approach, we defined the constant-roll condition so that we recover the USR model when constant-roll parameter approaches $-3$. Obviously, this condition is useful in studying the USR limit of a model. We dubbed these three conditions the FCR, HCR, and PCR conditions, respectively. These conditions were examined to determine under which conditions on action these purportedly distinct conditions result in homologous models and when they differ.
@@
To provide a clear example, we investigated a class of non-canonical models characterized by the constant sound speed, denoted as ”constant propagation inflation”, within the framework of our generalized constant-roll formalism. We demonstrated that within this class of models, the HCR and PCR conditions display homologous behavior. In a particular case with $c_{s}=1$ (corresponding to a class of models known as k/G inflation), we identified two distinct sets of constant-roll solutions under the HCR and FCR conditions. While these solutions generally differ, we showed that in specific limits where $H_{,\phi}$ approaches zero and the field settles at the extrema of the effective potential, both FCR and HCR models exhibit similar quasi-canonical behavior, where slow-roll parameters mimic their canonical counterparts, though the action generally deviates from the canonical form. We also conducted a full dynamical analysis of the constant-roll parameter under small perturbations in the field velocity to assess the stability of our exemplar of constant propagation model. We determined that for a range of the constant-roll parameters the analytic solutions are unstable and will generally transition to their corresponding dual attractor constant-roll solutions. We noted that the transition trajectory is universal in FCR case but is contingent upon the form of the action in HCR scenario.
@@
We also investigated another non-canonical action within the framework of the constant-roll scenario. We observed that when constant-roll parameter is negative, the model transitions automatically from a quasi-canonical phase to a canonical constant-roll phase, and when constant-roll parameter is positive it evolves in the opposite direction. We noted that while the nature of these two phases remains independent of the chosen constant-roll condition, the transition trajectory is highly dependent on this choice. We subsequently introduced a multi-stage ”constant $k$” scenario and conducted a numerical assessment of the system parameters and the power spectrum. This exploration demonstrated how modifying the constant-roll condition can significantly impact the model’s outcomes, emphasizing on the sensitivity of the system to the chosen constant-roll condition.
@@
We stress that the constant-roll conditions we defined in this work are not unique. Mathematically, one can define an infinite number of constant-roll conditions for a non-canonical model, each leading to the conventional constant-roll condition $\ddot{\phi}=\beta H\dot{\phi}$ in the canonical limit. We believe this is an intriguing flexibility of the model that has not yet been adequately examined. The study of a non-canonical model under the conventional constant-roll condition can indeed provide some insights. However, other forms of the constant-roll condition may also provide valuable information that should not overlooked.
@@ Appendix A The Turning Point of $n_{s}$ and $f_{NL}^{re}$
In this appendix we go into the mathematical details of the power spectrum and the contribution of field redefinition to non-Gaussianity in the constant-roll scenario. We explain why perturbations suddenly change behavior at $\beta=-3/2$. We begin by considering the general action (3.1),
@@
In this model, the Mukhanov-Sasaki variable $v_{k}$, which is related to the mode function of curvature perturbations $u_{k}$ through the relation $v_{k}\equiv zu_{k}$, satisfies the following equation [75, 78, 79]:
@@
$\frac{z^{\prime\prime}}{z}=a^{2}H^{2}\left(2-\epsilon_{1}+\frac{3}{2}\epsilon_{2}+\frac{1}{4}\epsilon_{2}^{2}-\frac{1}{2}\epsilon_{1}\epsilon_{2}+\frac{1}{2}\epsilon_{2}\epsilon_{3}+3q_{1}-\epsilon_{1}q_{1}+\epsilon_{2}q_{1}+q_{1}^{2}+q_{1}q_{2}\right),$ (A.3)
@@
At supper Hubble scales, $k\ll aH$, we can neglect the $c_{s}^{2}k^{2}$ term in equation (A.2) and integrated it out to find
$u=C_{1}+C_{2}\int\frac{\mathrm{~{}d}\tau}{a^{2}\epsilon_{1}}=C_{1}+\frac{C_{2}}{a_{0}^{3}H}\int\frac{\mathrm{~{}d}N}{e^{3N}\epsilon_{1}}.$ (A.5)
This approximate form is useful for analyzing the evolution of modes in large scales.
@@
By integrating this equation we find the slow-roll parameters plotted in Figure 1. However, since we are interested in the behavior of perturbations around $\beta_{\text{\tiny{H}}}=-3/2$ point, we will only consider solution (2.13). To solve equation (A.2), it is also necessary to derive $q_{j}$ parameters (A.4), which cannot be obtained without knowing the form of the action. For simplicity we consider the constant propagation model (4.1), where the sound speed is constant. Therefore, $q_{j}$ parameters are zero, and slow-roll parameters are approximated as $\epsilon_{1}\approx 0$, $\epsilon_{2}\approx 2\beta_{\text{\tiny{H}}}$, and $\epsilon_{3}\approx 0$ in large positive $N$. Using these results equation (A.2) simplifies to
@@
$\mathcal{P}_{s}(k)\equiv\frac{k^{3}}{2\pi^{2}}\left|u_{k}\right|^{2}=\frac{H^{2}}{8\pi^{2}\epsilon_{1}}\left(\frac{k}{aH}\right)^{3}\frac{\pi}{2}\left|H_{\nu}^{(1)}(-c_{s}k\tau)\right|^{2}.$ (A.11)
@@
$\mathcal{P}_{s}(k)\approx\frac{H^{2}}{8\pi^{2}\epsilon_{1}}\frac{2^{2\nu-1}[\Gamma(\nu)]^{2}}{\pi}\left(\frac{c_{s}k}{aH}\right)^{3-2\nu}.$ (A.12)
Accordingly, the spectral index is $n_{s}-1=3-2\nu=3-2\left|\beta_{\text{\tiny{H}}}+3/2\right|$, which is consistent with equation (2.35).
Our next step will be the evaluation of the contribution of field redefinition to non-Gaussianity around $\beta_{\text{\tiny{H}}}=-3/2$ in order to investigate its sudden behavior change. In the calculations of non-Gaussianity it is shown that a field redefinition as $\zeta\rightarrow\zeta_{n}+f(\zeta_{n})$ can significantly simplifies the interaction Hamiltonian. In general, for a $P(X)$ model (A.1), $f(\zeta)$ is given by
@@
If $\beta_{\text{\tiny{H}}}>-3/2$, modes will freeze out on super Hubble scales. Therefore, all derivative terms in above equation will vanish and will not contribute to non-Gaussianity. If $\beta_{\text{\tiny{H}}}<-3/2$, however, $\dot{\zeta}$ term is non-vanishing and will contribute. Therefore, for the calculation of the correlation function we only consider the first two terms in equation (A.13). After field redefinition $\zeta\rightarrow\zeta_{n}+\frac{\epsilon_{2}}{4c_{s}^{2}}\zeta_{n}^{2}+\frac{1}{c_{s}^{2}H}\zeta_{n}\dot{\zeta_{n}}$ the three point function becomes
@@
where $\mathcal{O}\left(\left\langle\zeta_{n}^{5}(\mathrm{x})\right\rangle\right)$ indicates the higher order terms. Writing the above expression in Fourier space and using relation $\langle{\zeta}_{n}\left(k_{1}\right)\zeta_{n}\left(k_{2}\right)\rangle=(2\pi)^{3}\delta\left(k_{1}+k_{2}\right)u_{k_{1}}u_{k_{2}}^{*}$, for the contribution of field redefinition to three point function we find
$(2\pi)^{3}\delta^{(3)}\left(\mathrm{x}_{1}+\mathrm{x}_{2}+\mathrm{x}_{3}\right)\left[\frac{\epsilon_{2}}{2c_{s}^{2}}\right.$ $\left.\left(\left|u\left(\mathrm{x}_{2}\right)\right|^{2}\left|u\left(\mathrm{x}_{3}\right)\right|^{2}+\mathrm{sym}\right)+\right.$ (A.15)
$\left.\frac{1}{c_{s}^{2}H}\right.$ $\left.\left(\left|u\left(\mathrm{x}_{2}\right)\right|^{2}\dot{u}\left(\mathrm{x}_{3}\right)u^{*}\left(\mathrm{x}_{3}\right)+\left|u\left(\mathrm{x}_{3}\right)\right|^{2}\dot{u}\left(\mathrm{x}_{2}\right)u^{*}\left(\mathrm{x}_{2}\right)+\mathrm{sym}\right)\right].$
We now try to find a relationship between $\dot{u}$ and $u$. For this, we use relation $v=zu$ to write
$\dot{u}=-uH\left(\frac{\tau v^{\prime}}{v}+\frac{\epsilon_{2}}{2}+1\right).$ (A.16)
@@
where again we have used asymptotic relation $\lim_{x\rightarrow 0}H_{\nu}^{(1)}(x)\simeq-\frac{i}{\pi}\Gamma(\nu)\left(x/2\right)^{-\nu}$ for Hankel function. In the light of equations (A.15), (A.16), and (A.17) the bispectrum and non-Gaussianity parameters can be found as

### v3

@@ 5.2 Evolution of Scalar Perturbations
$c_{s}^{2}$
$q_{1}$
$N$$N$$N$$N$$N$(a)(b)(c)(d)(e)CanonicFCRHCRPCR
Figure 11: The evolution of background parameters for the multiple-stage constant-roll scenario understudy in section (5.2). Panel: (a) constant-roll parameter $\beta$, (b) first slow-roll parameter $\epsilon_{1}$, (c) second slow-roll parameter $\epsilon_{2}$, (d) sound speed $c_{s}$, and (e) parameter $q_{1}$. In all panels (except panel (a)) the colour coding is the same. The green, blue, and yellow lines represent FCR, HCR, and PCR conditions, respectively. As is usual in comparative studies, we proposed the $\beta(N)$ function (in panel (a)) as a null hypothesis that makes no difference between models. There are four constant-roll stages with $\beta(N)\approx-0.017$, $-1.45$, $0.018$, and $1$, separated by three vertical grid lines at $N=10$, $20.5$, and $50$ $e$-folds. As shown in this figure, at early times the dynamics are in $\gamma X\gg 1$ regime (with $c_{s}^{2}=1/3$), but as $X$ decreases, it transitions to a stage in canonical regime. Our comparative study on the constant-roll models described in this section shows that the general behavior is similar but the transition period and its starting time is model dependent. In the PCR model, for example, the field spends more time in the non-canonical regime. The details of the scenario and point-by-point analysis of this figure are given in the text.
To illustrate a viable model in which a transient non-canonical constant-roll stage can generate the growth in the primordial curvature power spectrum, while satisfying CMB constraints and providing a graceful exit from inflation, we construct a four-stage constant-roll model. We plan to compare the effect of each constant-roll condition on the model; so we define constant-roll parameter as a function of time, $\beta(N)$, and use it in the three constant-roll conditions in a similar fashion. We define the following parameter:{sequation}[] β(N) =-495 (4241+tanh(3 (N-15))) (3940-tanh(3 (N-20.5))) (2827 + tanh(N-50)). This parameter (plotted in panel (a) of Figure 11) describes four stages of constant-roll inflation (with $\beta=-0.017$, $-1.45$, $0.018$, and $1$, respectively). The stages are connected using proper smooth functions. The constant-roll parameter is set to $-0.017$ in the first stage in order to comply with Planck data (see discussion after equation (2.36)). At this stage we normalized the power spectrum to meet the CMB constraints [76], $\mathcal{P}_{s}=2.1\times 10^{-9}$ at $k_{\text{cmb}}=0.05\text{Mpc}^{-1}$. After 15 $e$-folds we switch to a large $|\beta|$ stage responsible for enhancing the power spectrum. We take the constant-roll parameter smaller than $-1$ to have negative $z^{\prime}/z$, but larger than $-3/2$ in order to preserve the attractor behavior of the model. At $21.5$ $e$-folds, when the enhancement in the spectrum amplitude between small and large CMB scales reaches to $10^{6}-10^{7}$, we end this stage. The third stage is another slow-roll stage with $\beta=0.018$ that is compatible with Planck data (see discussion after equation (2.35)). At the end, we enter a short stage with large positive $\beta$ at 50 $e$-folds, that leads to an increase in $\dot{\phi}$ and accordingly $\epsilon_{1}$, allowing us to reach $\epsilon_{1}=1$ and exit inflation gracefully.
First, we use parameter (5.2) in equation (2.6) and solve the equations of motion and Mukhanov-Sasaki equation (A.2) numerically to derive power spectrum and dynamical parameters for canonical case. The results are illustrated in Figures 11 and 12 (dashed gray lines). As can be seen, the power has increased by about $10^{7}$ times, which is appropriate for PBH formation as suggested in references [67, 58]. In this scenario, as $\beta$ changes, the automatic switch to the corresponding solution will be needed. The first two stages can be described respectively by solutions (2.15) and (2.13), and solution (2.14) can be used for the last two stages.
For the non-canonical model (5.1) the situation is more involved. We repeat the above calculation for constant-roll conditions (5.9)-(5.11). Figures 11 and 12 show the results (the green, blue, and yellow lines represent FCR, HCR, and PCR conditions, respectively). It should be noted that here we want to compare three constant-roll conditions within a single scenario. We are concerned about the effect of changing the constant-roll condition on the model, not the compatibility of each power spectrum with the data. In this model, we start inflation in $\gamma X\gg 1$ phase. In the first stage, since $\beta$ is small and $\dot{\phi}$ decays slowly, there is not enough time to transit to the canonical case. During the second stage, $|\beta|$ increases significantly leading to a quick decay in $\dot{\phi}$ and the transition to the canonical case. The transition can be seen even more transparent in $c_{s}^{2}$ plot (panel (d) of Figure 11). It is evident that the model gradually transitions from a $\gamma X\gg 1$ phase with $c_{s}^{2}=1/3$ to the canonical phase with $c_{s}^{2}=1$. During the third stage, $\beta$ becomes positive and $\dot{\phi}$ begins to grow. However, since $\beta$ is small the system stays in the canonical regime throughout. In the last stage, as $\beta$ is large and positive, $\dot{\phi}$ grows dramatically fast and we transit back to $\gamma X\gg 1$ limit before exiting inflation.
Looking at panels (b) and (c) of Figure 11, it is evident that the slow-roll parameters for the HCR condition are identical to those for the canonical case. It is due to the fact that for the HCR condition slow-roll parameters can be derived irrespective of action. However, the power spectrum will be different, since the sound speed varies with time. As can be seen in left panel of Figure 12, the curve $z^{\prime}/(aHz)$ for the HCR model deviates slightly from the canonical case, which is the direct result of varying sound speed (see equations (5.21) and (A.3)).
Panel (b) of Figure 11 also shows that in all stages of the scenario except the third stage, the slope of $Log(\epsilon_{1})$ differs for three constant-roll conditions. This is because at third stage we reach the canonical case regardless of our choice of constant-roll condition. This point becomes more clear when we look at the second slow-roll parameter in panel (c) of Figure 11. Despite reaching a single path at about 20 $e$-folds, $\epsilon_{2}$ functions are different in earlier times when we have not reached the canonical limit yet. In order to get similar results in early times of inflation when $\gamma X$ is much larger than one we must choose $6\beta_{\text{\tiny{F}}}=3\beta_{\text{\tiny{H}}}=2\beta_{\text{\tiny{V}}}$ (see discussion after equation (5.11)), but in the scenario we utilized $\beta_{\text{\tiny{F}}}=\beta_{\text{\tiny{H}}}=\beta_{\text{\tiny{V}}}=\beta(N)$, where $\beta(N)$ is defined in (5.2). However, when we enter the canonical phase, our choice $\beta_{\text{\tiny{F}}}=\beta_{\text{\tiny{H}}}=\beta_{\text{\tiny{V}}}=\beta(N)$ leads to similar results for three constant-roll conditions.
Refer to caption      Refer to caption
@@
$Log(\mathcal{P}_{s})$
$N$$Log(k)$CanonicFCRHCRPCR
Figure 12: Parameter $z^{\prime}/(aHz)$ (left panel) and power spectrum $\mathcal{P}_{s}$ (right panel) for the four-stage constant $k$ scenario described by constant-roll parameter given in (5.2). The green, blue, and yellow lines represent the evolution constrained by FCR, HCR, and PCR conditions, respectively, and dashed gray curve corresponds to the constant role canonical model.
@@
Looking at the left panel of Figure 12, we see that in all cases $z^{\prime}/(aHz)$ transiently changes sign, which is necessary to enhance the power spectrum. The more negative it is, the stronger the driving force will enhance the power. Furthermore, we see that the transition time to the canonical phase is different for the three constant-roll conditions; in FCR case the model transitions to canonical case earlier while in PCR case this happens later. That means in FCR case some modes with longer wavelengths (compared to HCR and PCR cases) will also be affected by the enhancement mechanism. Therefore, by looking at $z^{\prime}/(aHz)$ we expect that the FCR condition will result in greater power spectrum enhancement. This can be seen in the right panel of Figure 11.
@@ 6 Summary and Conclusion
This study explored three methods for imposing the constant-roll condition on non-canonical models. We used a general action as a toy model to introduce a trio of constant-roll conditions, which are all equivalent in the canonical limit but may lead to different results in non-canonical cases. In the first approach, in which a more conventional form of the constant-roll condition in the literature is used, the condition was defined to ensure that the inflaton rolling rate remains constant throughout the inflation. Imposing another type of the constant-roll condition ensures a constant $\epsilon_{2}$ in the model. This second condition enabled us to derive the time profile of the Hubble and slow-roll parameters analytically, even without knowing the exact form of the action. For models with constant sound speed the power spectrum and spectral index of the model were also derived. In these two approaches, we may not obtain a flat USR potential for any value of the constant-roll parameter. In our third approach, we managed to acquire the USR model as constant-roll parameter approaches $-3$. We dubbed these three conditions the FCR, HCR, and PCR conditions, respectively. These conditions were examined to determine under which conditions on action these distinct conditions result in homologous models and when they differ.
@@
As an representative example in our comparative study on the generalized constant-roll models, we investigated a class of non-canonical models characterized by constant sound speed, denoted as ”constant propagation inflation”. We demonstrated that within this class of models, the HCR and PCR conditions display homologous behavior. In a particular case with $c_{s}=1$ (corresponding to a class of models known as k/G inflation), we identified two distinct sets of constant-roll solutions under the HCR and FCR conditions. While these solutions generally differ, we showed that in specific limits, where $H_{,\phi}$ approaches zero and the field settles at the extrema of the effective potential, both FCR and HCR models exhibit similar quasi-canonical behavior, where slow-roll parameters mimic their canonical counterparts, though the action generally deviates from the canonical form. We also conducted a full dynamical analysis of the constant-roll parameter under small perturbations in the field velocity to assess the stability of this exemplar of constant propagation model. We determined that for a range of the constant-roll parameters the analytic solutions are unstable and will generally transition to their corresponding dual attractor constant-roll solutions. We noted that the transition trajectory is universal in FCR case but is contingent upon the form of the action in HCR scenario.
@@
We also investigated a kinetically driven non-canonical model with varying sound speed. We observed that when constant-roll parameter is negative, the model transitions automatically from a non-canonical phase to a canonical constant-roll phase, and when constant-roll parameter is positive it evolves in the opposite direction. We noted that while the nature of these two phases remains independent of the chosen constant-roll condition, the transition trajectory is highly affected by this choice. We subsequently introduced a multi-stage constant-roll scenario and conducted a numerical assessment of the background parameters and the power spectrum of the curvature perturbations. This exploration also demonstrated the implications of model, emphasizing on the sensitivity of the system to the chosen constant-roll condition.
@@
We stress that the constant-roll conditions we defined in this work are not unique. Mathematically, one can define an infinite number of constant-roll conditions for a non-canonical model, each leading to the conventional constant-roll condition $\ddot{\phi}=\beta H\dot{\phi}$ in the canonical limit. We believe this is an intriguing flexibility of the model that has not yet been adequately examined. The study of a non-canonical model under the conventional constant-roll condition can indeed provide some insights. We believe that other forms of the constant-roll condition may also provide valuable information that should not overlooked; so avoiding those in favor of the conventional one is unreasonable.
@@ Appendix A The Turning Point of $n_{s}$ and $f_{NL}^{re}$
In this appendix, we go into the mathematical details of the power spectrum and non-Gaussianity in constant-roll scenarios. We specifically explain why in models with constant sound speed the perturbations’ behavior changes suddenly at $\beta=-3/2$. We begin by considering the general action (3.1),
@@
In this model, the Mukhanov-Sasaki variable $v_{k}$, which is related to the mode function of curvature perturbations $\mathcal{R}_{k}$ through the relation $v_{k}\equiv z\mathcal{R}_{k}$, satisfies the following equation [77, 80, 81]:
@@
$\frac{z^{\prime\prime}}{z}=a^{2}H^{2}\left(2-\epsilon_{1}+\frac{3}{2}\epsilon_{2}+\frac{1}{4}\epsilon_{2}^{2}-\frac{1}{2}\epsilon_{1}\epsilon_{2}+\frac{1}{2}\epsilon_{2}\epsilon_{3}-3q_{1}+\epsilon_{1}q_{1}-\epsilon_{2}q_{1}+q_{1}^{2}-q_{1}q_{2}\right),$ (A.3)
@@
At supper Hubble scales, $k\ll aH$, we can neglect the $c_{s}^{2}k^{2}$ term in equation (A.2) and the general solution for corresponding homogeneous curvature perturbation is given by
$\mathcal{R}=C_{1}+C_{2}\int\frac{\mathrm{~{}d}\tau}{z^{2}},$ (A.5)
where $C_{1}$ and $C_{2}$ are constants of integration. This approximate form is useful for analysing the evolution of modes in large scales and de Sitter space.
@@
By integrating this equation we find the slow-roll parameters plotted in Figure 1. However, since we are interested in the behavior of perturbations around $\beta_{\text{\tiny{H}}}=-3/2$ point, we will only consider solution (2.13). To solve equation (A.2), it is also necessary to derive $q_{j}$ parameters, which cannot be obtained without knowing the form of the action. For simplicity, we consider the constant propagation model (4.1), where the sound speed is constant. Therefore, $q_{j}$ parameters are zero, and slow-roll parameters are approximated as $\epsilon_{1}\approx 0$, $\epsilon_{2}\approx 2\beta_{\text{\tiny{H}}}$, and $\epsilon_{3}\approx 0$ in large positive $N$. Using these results equation (A.2) simplifies to
@@
$\mathcal{P}_{s}(k)\equiv\frac{k^{3}}{2\pi^{2}}\left|\mathcal{R}_{k}\right|^{2}=\frac{H^{2}}{8\pi^{2}\epsilon_{1}}\left(\frac{k}{aH}\right)^{3}\frac{\pi}{2}\left|H_{\nu}^{(1)}(-c_{s}k\tau)\right|^{2}.$ (A.11)
@@
$\mathcal{P}_{s}(k)\approx\frac{H^{2}}{8\pi^{2}\epsilon_{1}}\frac{2^{2\nu-1}[\Gamma(\nu)]^{2}}{\pi}\left(\frac{c_{s}k}{aH}\right)^{3-2\nu},$ (A.12)
which is basically the equation (44) of reference [13] with an arbitrary (but constant) sound speed. Accordingly, the spectral index is $n_{s}-1=3-2\nu=3-2\left|\beta_{\text{\tiny{H}}}+3/2\right|$, which is consistent with equation (2.35).
Our next step will be the evaluation of the contribution of field redefinition to non-Gaussianity around $\beta_{\text{\tiny{H}}}=-3/2$ in order to investigate its sudden behavior change. In the calculations of non-Gaussianity, it is shown that a field redefinition as $\zeta\rightarrow\zeta_{n}+f(\zeta_{n})$ can significantly simplify the interaction Hamiltonian. In general, for a $P(X,\phi)$ action (A.1), $f(\zeta)$ is given by
@@
If $\beta_{\text{\tiny{H}}}>-3/2$, modes will freeze out on super Hubble scales. Therefore, all derivative terms in the above equation vanish and do not contribute to the non-Gaussianity. If $\beta_{\text{\tiny{H}}}<-3/2$, however, $\dot{\zeta}$ term is non-vanishing and will contribute. Therefore, for the calculation of the correlation function we only consider the first two terms in equation (A.13). After field redefinition $\zeta\rightarrow\zeta_{n}+\frac{\epsilon_{2}}{4c_{s}^{2}}\zeta_{n}^{2}+\frac{1}{c_{s}^{2}H}\zeta_{n}\dot{\zeta_{n}}$ the three-point function becomes
@@
where $\mathcal{O}\left(\left\langle\zeta_{n}^{5}(\mathrm{x})\right\rangle\right)$ indicates the higher order terms. Writing the above expression in Fourier space and using relation $\langle{\zeta}_{n}\left(k_{1}\right)\zeta_{n}\left(k_{2}\right)\rangle=(2\pi)^{3}\delta\left(k_{1}+k_{2}\right)\mathcal{R}_{k_{1}}\mathcal{R}_{k_{2}}^{*}$, for the contribution of field redefinition to three-point function we find
$(2\pi)^{3}\delta^{(3)}\left(\mathrm{x}_{1}+\mathrm{x}_{2}+\mathrm{x}_{3}\right)\left[\frac{\epsilon_{2}}{2c_{s}^{2}}\right.$ $\left.\left(\left|\mathcal{R}\left(\mathrm{x}_{2}\right)\right|^{2}\left|\mathcal{R}\left(\mathrm{x}_{3}\right)\right|^{2}+\mathrm{sym}\right)+\right.$ (A.15)
$\left.\frac{1}{c_{s}^{2}H}\right.$ $\left.\left(\left|\mathcal{R}\left(\mathrm{x}_{2}\right)\right|^{2}\dot{\mathcal{R}}\left(\mathrm{x}_{3}\right)\mathcal{R}^{*}\left(\mathrm{x}_{3}\right)+\left|\mathcal{R}\left(\mathrm{x}_{3}\right)\right|^{2}\dot{\mathcal{R}}\left(\mathrm{x}_{2}\right)\mathcal{R}^{*}\left(\mathrm{x}_{2}\right)+\mathrm{sym}\right)\right].$
We now try to find a relationship between $\dot{\mathcal{R}}$ and $\mathcal{R}$. For this, we use relation $v=z\mathcal{R}$ to write
˙$\dot{\mathcal{R}}=-\mathcal{R}H\left(\frac{\tau v^{\prime}}{v}+\frac{\epsilon_{2}}{2}+1\right).$ (A.16)
@@
where again we have used asymptotic relation $\lim_{x\rightarrow 0}H_{\nu}^{(1)}(x)\simeq-\frac{i}{\pi}\Gamma(\nu)\left(x/2\right)^{-\nu}$ for Hankel function. In the light of equations (A.15)-(A.17) the bispectrum and non-Gaussianity parameter can be found as